OpenStudy (anonymous):

IF a vector F=2i+4j+8k is placed at (2,2,2) What will be angle with y axis to this vector.?

OpenStudy (anonymous):

Sine is the trig function that gives the Y value, and the j unit vector also corresponds to the Y, so using the triangle formed between the Y-axis and the vector... Use sine^-1 theta to solve for the angle.

OpenStudy (anonymous):

|F|=sqrt(84) \[\theta_y=\cos^{-1}(\frac{F.j}{|F|})\]

OpenStudy (anonymous):

Does that answer your question?